OpenStudy (anonymous):

What is the sum of all positive integers r that satisfy Lcm[r,700] = 7000

ganeshie8 (ganeshie8):

Hint : \[ab = \text{lcm}(a,b) \times \gcd(a,b) \]

OpenStudy (anonymous):

8000 (1000, 7000)

ganeshie8 (ganeshie8):

Yep! how did you get that ?

OpenStudy (mathmath333):

\(\)

OpenStudy (anonymous):

Note the prime factorizations $700=2^2 * 5^2 * 7 and 7000=2^3 * 5^3 * 7 1) 2^3 * 5^3 = 1000 2) 7*2^3*5^3= 7000